### v1

# Enhancing operational wind downscaling capabilities over Canada: Application of a Conditional Wasserstein GAN methodology

Abstract
Wind downscaling is essential for improving the spatial resolution of weather forecasts, particularly in operational Numerical Weather Prediction (NWP). This study advances wind downscaling by extending the DownGAN framework introduced by Annau et al. to operational datasets from the Global Deterministic Prediction System (GDPS) and High-Resolution Deterministic Prediction System (HRDPS), covering the entire Canadian domain. We enhance the model by incorporating high-resolution static covariates, such as HRDPS-derived topography, into a Conditional Wasserstein Generative Adversarial Network with Gradient Penalty, implemented using a UNET-based generator. Following the DownGAN framework, our methodology integrates low-resolution GDPS forecasts (15 km, 10-day horizon) and high-resolution HRDPS forecasts (2.5 km, 48-hour horizon) with Frequency Separation techniques adapted from computer vision. Through robust training and inference over the Canadian region, we demonstrate the operational scalability of our approach, achieving significant improvements in wind downscaling accuracy. Statistical validation highlights reductions in root mean square error (RMSE) and log spectral distance (LSD) metrics compared to the original DownGAN. High-resolution conditioning covariates and Frequency Separation strategies prove instrumental in enhancing model performance. This work underscores the potential for extending high-resolution wind forecasts beyond the 48-hour horizon, bridging the gap to the 10-day low resolution global forecast window.
1 Introduction
Accurate high-resolution wind forecasts are crucial for several application sectors such as aviation [1], wind energy [2], power infrastructure [3], public safety [4], wildfire management [5], avalanche warnings [6], winter sports [7], air quality [8] and water quality management [9]. Typically, wind forecasts are obtained from numerical weather predictions (NWPs), starting with a global forecast model followed by a limited-area forecast model which takes the lower-resolution model as boundary conditions in a procedure called dynamical downscaling. In Canada, high-resolution numerical prediction models have a limited forecast horizon of 48 hours due to computational constraints. An applied perspective of the work presented in this paper is to extend high-resolution wind forecast services from 48 hours to the same 10 days forecast horizon as the Canadian global forecast model.
In meteorology, downscaling is the operation of passing from a coarse (or low-resolution) field to a fine (or high-resolution) field. Besides dynamical downscaling, statistical downscaling is another well established approach for correcting wind at station locations. Several variations of Model Output Statistics (MOS) [10, 11] have been proposed in the literature, with the Updateable Model Output Statistics (UMOS) [12] approach the one implemented operationally at Environment and climate Change Canada (ECCC). All the MOS techniques are essentially based on the concept of Multiple Linear Regression (MLR), where several predictors from the lower-resolution NWP output are combined linearly to minimize the mean square error against an observed weather field, with one specific model used at each station location. MOS techniques can be expanded to gridded fields as long as a high-resolution gridded target is available, or specific corrections applied at each station can be interpolated to the grid [13], but they cannot capture complex non-linear spatial interactions between variables.
More recently, several contributions have demonstrated the potential of Artificial Intelligence (AI) based downscaling [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25], applying successfully techniques originally developed for image processing and computer vision to weather data. In the field of image processing, the image super-resolution (or zooming) problem consists of taking a coarser image as input and producing a finer image at higher resolution. [^1] Confusingly, what is called upsampling (downsampling) in the field of image processing is called downscaling (upscaling) in meteorology.
Earlier methods were achieving an enhancement factor of 2 to 4 times, but with the advent of deep learning, an enhancement factor of 8 and more is now possible. Generative AI in particular has proven to be a powerful tool not only to artificially produce realistic-looking images and videos from a random seed, but also to condition the generative AI on coarser or partially missing data (called inpainting).
AI-based techniques for downscaling weather and climate fields can be classified in an increasing order of complexity and recency. At the end of spectrum, ensemble decision trees techniques such random forests and extreme gradient boosting are the most similar to MLR and the easiest to implement. Other classical machine learning techniques such as support vector machine and shallow artificial neural networks with a single hidden layer can also fit in that category. Methods based on Convolutional Neural Networks (CNNs) have proved to be both popular and powerful. Generative Adversarial Networks (GANs) are the first type of generative AI allowing to sample an ensemble of possible realizations [15]. More recently, latent diffusion techniques have emerged as an interesting alternative to GAN. Finally, methods based on vision transformers and in particular those that consist of fine-tuning a weather foundation model are emerging.
In the literature, these methods have been applied to a variety of weather fields such as temperature, wind speed or wind vectors, and precipitation. Many of the published work focus on one or several small areas as a proof-of-concept, and few are used in an operational setting. It is important to note that similar AI techniques can be used either for downscaling weather forecasts, climate projections or reanalysis data, with the main difference in the choice and resolution of source and target data. Moreover, with a general downscaling model , there is no conceptual difference between downscaling to a finer grid or downscaling to a point location that could be either an existing weather station or any location of interest.
The key contributions of this work are as follows: 1) We extend the DownGAN framework introduced by [15] by applying its methods to operational Numerical Weather Prediction (NWP) datasets, specifically GDPS and HRDPS across the entire Canadian domain, expanding beyond the regional focus of the original study. 2) We enhance the DownGAN model, a Wasserstein GAN with Gradient Penalty (WGAN-GP) framework for wind downscaling, by introducing high-resolution static covariates, such as topography from the HRDPS domain, as conditioning inputs within the generator architecture, implemented using a UNET. We refer this modified framework as a Conditional WGAN-GP. 3) We propose a robust training and inference schema that allows the model to effectively learn wind patterns and generate predictions over the entire Canadian region. This development ensures the scalability and operational relevance of the methodology. 4) We perform statistical validation for large-scale wind downscaling across the Canadian domain, identifying limitations such as the ineffectiveness of Annau’s partial frequency separation technique in this context. However, Frequency Separation methods adapted from computer vision and used by DownGAN show strong performance. The validation highlights the importance of high-resolution conditioning covariates and the utility of Frequency Separation in this context.
The paper is divided as follows. Section 2 introduces the datasets and the pre-processing steps in preparation for training the AI model. Section 3 presents the methods for training and inference of the AI model. Experimental results follow in Section 4, and Discussion and Conclusions are presented in Section 5.
2 Dataset
2.1 Global Deterministic Prediction System
The Global Deterministic Prediction System (GDPS) [26] is ECCC’s operational long-range forecast model. It represents the global configuration of the Global Environmental Multiscale (GEM) model [27] and it is run twice a day (at 00 UTC and 12 UTC), producing 10-day forecasts at a 15-km nominal resolution.
2.2 High-Resolution Deterministic Prediction System
The High-Resolution Deterministic Prediction System (HRDPS) [28] is a limited-area model (LAM) configuration of GEM for Canada and Northern United States. The prediction is run operationally at ECCC every 6 hours (00 UTC, 06 UTC, 12 UTC and 18 UTC runs) for a 48 hours forecast at 2.5-km nominal resolution. The HRDPS is the main source of NWP guidance for the Meteorological Service of Canada for day 1 and day 2 forecasts.
Figure 1: GDPS and HRDPS
2.3 Predictors and Predictands
High-resolution zonal (East-West) and meridional (North-South) surface (10 m) wind components from HRDPS are used as the target variables. By predicting both components of the wind vector, we can derive both wind speed and direction at the resolution of HRDPS. Low-resolution zonal and meridional surface wind from GDPS are taken as predictors along with surface temperature, as well as wind, temperature and vertical motion at 546 hPa. In contrast with [15], high-resolution geophysical fields are added as predictors. These static high-resolution covariates can be used as predictors since these variables are constant at the time-scale of weather prediction. The topography field is particularly important in mountainous regions where wind channeling is predominant. The land-water mask not only provides an indication of changing surface roughness between the smoother water and other land, but can also be a driver of local winds such as sea breezes and lake breezes. Finally, the surface roughness variable modulates the speed of the wind and changing surface roughness can drive turbulence in the planet boundary layer. Table LABEL:table:vars_description describes the meteorological variables used for training the AI downscaling model.
Table 1: Description of Variables for GDPS and HRDPS
Description
GDPS Predictors
True geographical West-East (zonal) component of the horizontal wind at the surface (10 m) [m/s].
True geographical South-North (meridional) component of the horizontal wind at the surface (10 m) [m/s].
Air temperature at the surface (1.5 m) [°C].
Air temperature vertically interpolated at 546 hPa [°C].
True geographical West-East (zonal) component of the horizontal wind vertically interpolated at 546 hPa [m/s].
True geographical South-North (meridional) component of the horizontal wind vertically interpolated at 546 hPa [m/s].
Vertical motion vertically interpolated at 546 hPa [Pa/s].
HRDPS Predictands
True geographical West-East (zonal) component of the horizontal wind at the surface (10 m) [m/s]. Denoted by u10.
True geographical South-North (meridional) component of the horizontal wind at the surface (10 m) [m/s]. Denoted by v10.
HRDPS Static Covariates
Model orography [m].
Water/land mask [fraction].
Roughness length [m].
2.4 Data Pre-processing
The map projection of the GDPS outputs is a Yin-Yang grid with a different rotated latitude-longitude map projection than the one of the HRDPS grid. We regrid the GDPS grid to the HRDPS grid using nearest neighbor interpolation. The interpolated data is then reduced by a factor of 8 (20-km nominal resolution) as the input of the AI downscaling method. Fig. 1 illustrates the respective domains of GDPS (cropped over North-America) and of HRDPS.
3 Methods
In this study, we adopt the deep learning framework for downscaling wind components proposed by [15]. This framework is based on Wasserstein Generative Adversarial Networks with Gradient Penalty (WGAN-GP) [29] and incorporates elements from Generative Adversarial Networks for image super-resolution (SRGAN) [30] as a baseline for the downscaling task. In the following sections, we elaborate on the application of WGANs for the downscaling problem, describe the model architecture, and introduce our approach to integrating high-resolution static covariates into the training process using UNETs [31].
3.1 Conditional WGAN-GP for Downscaling
Given a training set consisting of dynamic low-resolution (GDPS) and high-resolution (HRDPS) forecast hours, denoted as ${\bf x}$ and ${\bf y}$, respectively. A conditional WGAN-GP, conditioned on the covariates ${\bf y_{cov}}$ in the HRDPS domain, aims to minimize the following losses for the critic $\mathcal{L}_{C}$ and generator $\mathcal{L}_{G}$ [15, 29, 32]:
$\mathcal{L}_{C}=\mathbb{E}_{G({\bf x})\sim\mathbb{P}_{g}}[C(G({\bf x}\mid{\bf y_{cov}}))]-\mathbb{E}_{{\bf y}\sim\mathbb{P}_{r}}[C({\bf y})]+\lambda\mathbb{E}_{{\bf\hat{y}}\sim\mathbb{P}_{\hat{y}}}[(\|\nabla_{\hat{y}}C(\hat{y})\|_{2}-1)^{2}]$ (1)
$\mathcal{L}_{G}=-\gamma\mathbb{E}_{G({\bf x})\sim\mathbb{P}_{g}}[C(G({\bf x}\mid{\bf y_{cov}}))]+\alpha\mathbb{E}_{{\bf y}\sim\mathbb{P}_{r}}l_{c}({\bf y},G({\bf x}\mid{\bf y_{cov}}))$ (2)
In this formulation, $C$ denotes a real-valued function known as the Critic, which is defined on the HRDPS domain. The function $G$ represents the Generator, which maps low-resolution GDPS to the HRDPS domain. The symbol $\mathbb{P}_{g}$ denotes the conditional distribution of the generated downscaled GDPS data given the covariates ${\bf y_{cov}}$, while $\mathbb{P}_{r}$ represents the distribution of the HRDPS points. The expression $\mathbb{E}_{{\bf\hat{y}}\sim\mathbb{P}_{\hat{y}}}[(\|\nabla_{\hat{y}}C(\hat{y})\|_{2}-1)^{2}]$ represents the gradient penalty, which is incorporated into the formulation to ensure valid Wasserstein distances by enforcing 1-Lipschitz continuity for the Critic. Here, ${\bf\hat{y}}\sim\mathbb{P}_{\hat{y}}$ denotes a linearly interpolated generated value defined by ${\bf\hat{y}}=\epsilon{\bf y}+(1-\epsilon)G({\bf x}\mid{\bf y_{cov}})$, where $\epsilon\sim U[0,1]$. The scalars $\lambda$, $\gamma$, and $\alpha$ are regularization hyperparameters: $\lambda$ controls the gradient penalty term, $\gamma$ the adversarial loss term, $\mathbb{E}_{G({\bf x})\sim\mathbb{P}_{g}}[C(G({\bf x}\mid{\bf y_{cov}}))]$ and $\alpha$ the content loss term, $\mathbb{E}_{{\bf y}\sim\mathbb{P}_{r}}l_{c}({\bf y},G({\bf x}\mid{\bf y_{cov}}))$. All expectations in the equations above are computed across all dimensions of ${\bf x}$ and ${\bf y}$. Specifically, ${\bf x}$ is a tensor of size $B\times C_{l}\times W_{l}\times H_{l}$, and ${\bf y}$ is a tensor of size $B\times C_{h}\times W_{h}\times H_{h}$. Here, $B$ represents the batch size, $W$ and $H$ denote the spatial width and height, respectively, while $C$ indicates the number of input channels (variates). The subscripts $l$ and $h$ refer to the GDPS and HRDPS datasets, respectively. Additionally, ${\bf y_{cov}}$ is a tensor with dimensions $C_{cov}\times W_{l}\times H_{l}$, where values are replicated across the batch size $B$. This replication is due to the covariates being static and independent of time. It is important to note that if ${\bf y_{cov}}$ is not provided, the WGAN-GP model reduces to the formulation presented in [15].
Utilizing WGAN-GP for downscaling is preferred over traditional GANs due to the advantages offered by WGANs. These advantages include a more stable training process and the provision of meaningful loss metrics, which contribute to achieving better generative outputs. Additionally, WGAN-GP effectively mitigates common challenges encountered in GAN training, such as mode collapse and vanishing gradients. For a more comprehensive understanding of these concepts, please refer to the relevant literature [15, 29, 33].
3.1.1 Frequency Separation
The core idea of the frequency separation approach is to decompose the low- and high-frequency components and treat them differently during WGAN-GP training [15, 34], with the aim of enhancing the quality of downscaled wind components. This allows the adversarial training to focus specifically on generating the high-frequency components, which are crucial for capturing fine details. In this approach, the high-frequency components are derived as ${\bf y}_{h}={\bf y}-H({\bf y})$ for the HRDPS data, and $G({\bf x}\mid{\bf y_{cov}})_{h}=G({\bf x}\mid{\bf y_{cov}})-H(G({\bf x}\mid{\bf y_{cov}}))$ for the generated downscaled data, where $H$ represents a low-pass filter. In this case, during WGAN-GP training, the Critic will focus on optimizing the loss based only in the high frequency components ${\bf y}_{h}$ and $(G({\bf x}\mid{\bf y_{cov}})_{h}$ instead of ${\bf y}$ and $(G({\bf x}\mid{\bf y_{cov}}))$. On the other side the Generator delegates the generation of high frequency components to the adversarial component $-\mathbb{E}_{G({\bf x})_{h}\sim\mathbb{P}_{g}}[C(G({\bf x}\mid{\bf y_{cov}}))_{h}]$ and the low frequency components to the content loss $\mathbb{E}_{{\bf y}\sim\mathbb{P}_{r}}l_{c}(H({\bf y}),H(G({\bf x}\mid{\bf y_{cov}})))$. In [15] is proposed the partial frequency separation, this approach is similar than before but the adversarial loss is used in its original formulation, i.e., $-\mathbb{E}_{G({\bf x})\sim\mathbb{P}_{g}}[C(G({\bf x}\mid{\bf y_{cov}}))]$
3.2 Architectures
3.2.1 Critic Network
To implement the proposed conditional WGAN-GP, we employed the same Critic network described by [15], as illustrated in Figure 2. The figure provides a schematic of the Critic network, with the numbers above each block indicating the dimensions of the data representations at various stages. The labels below the blocks denote the kernel size $k$, the number of filters $n$, and the stride $s$ used in each layer. The Critic utilizes a VGG-style architecture that progressively reduces the spatial dimensions of the input, which may include HRDPS wind components or the downscaled version of GDPS wind components provided by the Generator network, through a sequence of convolutional layers. The network processes the input data using 3x3 convolutions with LeakyReLU activations, where each layer increases the number of feature maps while downsampling the spatial dimensions using a stride of 2. This approach allows the network to expand the number of channels up to eight times the input dimension. The extracted features are then passed through a final processing block, a regression-style scoring mechanism that reduces the feature vector to a single scalar output. This scalar score, used as an approximation of the Wasserstein distance, indicates the Critic’s assessment of how real the input data appears.
Figure 2: Schematic of the Critic Network
3.2.2 Generator Network
The generator network implemented in the proposed model draws upon concepts from [15, 30] and is structured according to a U-Net architecture. The architecture is depicted in Figure 3 We employ this design because it facilitates the integration of HRDPS covariates and GDPS variates as inputs. The Generator embeds the HRDPS covariates through a sequence of downsampling layers reducing the spatial dimensions by a factor of eight. The resulting output tensor is then concatenated with the feature map tensor derived from the GDPS variates. This combined tensor passes through a series of residual-in-residual dense blocks, for a deep feature extraction. Following this stage, the network employs a sequence of upsampling layers that utilize pixel shuffle operations, enabling the progressive restoration of spatial resolution. Throughout the upsampling process, skip connections from the downsampling layers are leveraged to facilitate the merging of local and global features while preserving high-frequency details. The final output layer incorporates a convolutional operation to generate the GDPS downscaled version of wind componets. This U-Net architecture serves as the mechanism for implementing the conditional Wasserstein GAN, effectively utilizing HRDPS covariates to guide the generative process GDPS wind components.
Figure 3: Schematic of the Generator Network
4 Experiments
4.1 Dataset Split
This section outlines the experimental setup, using a dataset of matched GDPS and HRDPS forecast hours.For each prediction system, 24 hours of forecasts are used for each day consisting of forecast lead-times 6-17 hours for each of the 00 and 12 UTC runs. In this way, we avoid both the model spin-up time, and later forecast times where the two systems might have diverged because of error growth. For training, we selected all GDPS and HRDPS files with predictions starting between July 2022 to June 2023. The dataset was randomly split into 7,150 files for training, 1,178 for validation, and 1,162 for testing.
4.2 Metrics
To evaluate the performance of the proposed conditional WGAN-GP, we used the Root Mean Squared Error (RMSE) between the HRDPS reference and the downscaled GDPS wind components. Additionally, we employed the Radially Averaged Log-Spectrum Distance (LSD) [19]:
$LSD(\overline{P}_{true},\overline{P}_{pred})=\sqrt{\frac{1}{N_{r}}\sum_{r}\left(10\log_{10}\frac{\overline{P}_{true}(r)}{\overline{P}_{pred}(r)}\right)^{2}},$ (3)
where $P(r)$ is the Radially Averaged Power Spectrum Density [19] Observe that LSD quantifies the difference between predicted and true power spectra on a logarithmic scale. Subscripts pred and true are the $P(r)$ of predicted and true wind fields, respectively.
4.3 Training Procedures
The GDPS and HRDPS datasets cover the entire Canadian region. We applied a nearest-neighbor regridding approach to align the GDPS data with the HRDPS grid of $(2540,1280)$ spatial values, because each data set was provided in different rotated coordinate systems [^2] The regridding process was done using the xESMF Python Library [35]. We use this regridding because preserves original pixel values avoiding interpolation artifacts. In this sense, each GDPS and HRDPS file is represented by tensors of sizes$(7,2540,1280)$ and $(2,2540,1280)$, where 7 and 2 represent the variable counts in GDPS and HRDPS, respectively, as detailed in Table LABEL:table:vars_description. To train the WGAN-GP, we based our implementation in the open-source implementation from [36]. We used the following hyperparameters: critic Iterations: 5 (indicating one generator update per five critic updates), batch size: 32, learning Rate: 0.00025, regularization hyperparameters: $\lambda=10$ (gradient penalty), $\gamma=0.01$ (adversarial loss), and $\alpha=5$ (content loss). We used spatial random cropping for training with a $(128\times 128)$ crop size and a downsampling factor of 8 for the regridded GDPS, resulting in GDPS tensors of $(7,16,16)$ for each corresponding HRDPS tensor of $(2,128,128)$. We generated 832 crops per GDPS/HRDPS pair, amounting to 5,948,800 random crops for the 7,150 pairs in the training dataset. Here, one ”epoch” corresponds to using all 832 random crops from a single GDPS-HRDPS forecast hour. This random crop setup allows six generator updates using a batch size of 32.
The UNET generator network is configured with inputs of $(32,7,16,16)$ (batch size, variables, spatial dimensions) and receives an HRDPS static covariate tensor of $(3,128,128)$. The generator produces tensors of $(32,2,128,128)$, representing the downscaled u10 and v10 wind components in the HRDPS domain. The critic network processes input tensors of $(32,2,128,128)$, derived from either HRDPS data or the downscaled GDPS output. The proposed conditional WGAN-GP was trained for 65,250 epochs, with validation involving 32 random crops per GDPS-HRDPS file in the validation set.
4.4 Effect of using HRDPS Static Covariates
In this experiment, we investigate the impact of incorporating HRDPS static covariates on downscaling accuracy. To this end, we compare the DownGAN model [15] using the open-source implementation from [36], against a the proposed conditional WGAN-GP that utilizes HRDPS covariates. Both models are trained with identical hyperparameters and configurations to ensure a fair comparison. The primary difference lies in their generator architectures: DownGAN employs a generator based on the SRGAN network introduced by [30], while the WGAN-GP model adapts the DownGAN generator to a UNET structure, allowing it to effectively incorporate the additional HRDPS covariates.
Figure 4: Training and validation curves: DownGAN vs conditional WGAN-GP
Figure 4 presents the MSE curve trajectories for the DownGAN and the Conditional WGAN-GP models across training and validation epochs. The MSE values are averaged over intervals of 7250 epochs, which correspond to the number of forecast hours in the training set. In the plot, the DownGAN model’s training and validation MSE average values are represented by blue and red markers, respectively, while the Conditional WGAN-GP model’s MSE average values are shown in green (training) and orange (validation). The Conditional WGAN-GP model, which incorporates HRDPS static covariates, demonstrates consistently lower MSE values compared to the DownGAN model. The minimum validation MSE average value for each model is marked with an ”x”. The Conditional WGAN-GP model attains a lower validation error across all intervals, suggesting its enhanced capability for downscaling accuracy.
4.5 Effect of Frequency Separation
This section shows the experimental results to investigate the impact of Frequency Separation (FS) and Partial Frequency Separation (PFS) on the performance of the conditional WGAN-GP model. We tested multiple configurations for both FS and PFS, varying the low-pass average filter kernel sizes. Specifically, we used kernel sizes of $(5,5),(9,9)$ and $(13,13)$ for each WGAN-GP model with FS and PFS. Additionally, we included a conditional WGAN-GP model without FS or PFS as a baseline (i.e., same MSE validation curve from Figure 4). For the experiments, we applied transfer learning starting from epoch 21,750, and continuing training up to epoch 50,750. Figure 5 presents a comparison of the models’ performance based on the MSE validation metric.
Figure 5: Effect of Frequency separation in the validation curve
From the experiments, we observed that FS is in general better than PFS, and than small filter kernel sizes are better in term of the validation MSE metric. For this experiment, WGAN-GP with FS achieved the best configuration using a filter size of $(5,5)$, obtaining an average MSE validation score of 0.86 in the interval $[43500,50750]$. This result outperformed both the best PFS configuration (filter size of $(5,5)$ and MSE validation score of 0.93) and the case without frequency separation (MSE validation score of 0.97)
Figure 6: Effect of Frequency separation in downscaling at epoch 36250 — Left to right images: GDPS, GDPS downscaled with no frequency separation case, GDPS downscaled with WGAN-GP with FS using filter $(7,7)$ and HRDPS reference
Figure 6 shows the effect of FS in dealing with high frequencies as a function of epochs 29000, 36250, 43500 and 50750. In this case, checkboard artifacts are less evident in WGAN-GP with FS than in the case with no frequency separation.
4.6 Metrics on the Test Set
This section presents the metrics computed on a test set comprising 1,162 GDPS forecast hours. Figure 7 displays the results, contrasting the 1,162 downscaled GDPS forecast hours against HRDPS references in terms of RMSE and LSD metrics. We used Bilinear interpolation and nearest neighbor (NN) interpolation as baseline methods. Additionally, we evaluated several models, including the DownGAN model at checkpoint 29000, the Conditional WGAN-GP model at checkpoints 21,750 and 43,500 (no frequency separation scenario: NFS - 21,750 and NFS - 43,500, respectively), and the Conditional WGAN-GP model with frequency separation with kernels $(13,13),(9,9)$ and $(5,5)$ at checkpoints 50,750 denoted by FS13 - 50,750, FS9 - 50,750, and FS5 - 50,750 respectively. The decision to select a specific model checkpoint was based on the lowest validation MSE metric (See experiments from Section 4.5).
Figure 7: RMSE and LSD Metrics on the Test Set
As shown by Figure 7, based on the test error in both metrics for the two wind components, the interpolation methods performed the worst. Additionally, the DownGAN model exhibited worse performance than the Conditional WGAN-GP variations. Among the Conditional WGAN-GP models, the method with frequency separation and a kernel size of $5,5$ provided the best performance in terms of the RMSE metric, and the one with kernel size of $9,9$ in terms of the LSD metric Table LABEL:tab:metrics_summary presents a summary of the metrics on the test set in terms of their mean and standard deviation statistics.
Table 2: Summary of Results.
Method RMSE (u10) RMSE (v10) LSD (u10) LSD (v10)
Mean Std Mean Std Mean Std Mean Std
DownGAN - 30000 0.953 0.333 0.904 0.305 2.287 1.223 2.327 1.292
FS13 - 50750 0.785 0.296 0.785 0.293 1.820 0.939 1.852 0.980
FS9 - 50750 0.771 0.298 0.764 0.286 1.755 0.914 1.671 0.887
FS5 - 50750 0.749 0.296 0.744 0.287 1.842 0.964 1.756 0.934
NFS - 43500 0.803 0.311 0.778 0.294 1.889 0.973 1.847 0.970
NFS - 21750 0.843 0.326 0.812 0.303 1.962 0.992 2.082 1.078
Bilinear 1.012 0.396 0.993 0.378 3.462 2.379 3.476 2.412
KNN 1.041 0.399 1.023 0.383 3.994 0.790 4.008 0.764
4.7 Inference over the whole Canadian Region
Downscaling is crucial for analyzing large geographical regions; however, memory constraints can arise due to factors such as spatial resolution, data quantity, and model size. To address this issue, downscaling is often performed on smaller regions or patches, which are then stitched together to create a larger image. To downscale wind components on the entire Canadian domain we used the conditional WGAN-GP to generate overlapping $128\times 128$ images across the domain and then blend these overlapping areas using a weight matrix $W$, which is predefined $128\times 128$ matrix with weights in the interval $[0,1]$, increasing radially toward the center. Specifically, to generate a larger downscaling image $I$ from smaller $128\times 128$ images (patches), the algorithm starts by generating consecutive non-overlapping patches with a stride of 128 and copying the result directly into the corresponding region of $I$. Then, the entire region is again parsed, but this time the new patches overlap the artefacts. The final image is created by using linear interpolation between the patches generated in the last step and the correspondent pixels that were already in $I$ by using the weight matrix $W$. Each value for pixel $ij$ in image I’ is given by $I_{ij}^{\prime}=I_{ij}*(1-W_{ij})+\text{patch}_{ij}*W_{ij}$ Figure 8 provides a visual representation of this approach.
Figure 8: Inference overlapping strategy used to downscaling on the whole Canadian Region
The top part of the figure highlights the region $I$, which contains artifacts and is constructed using consecutive non-overlapping patches. The bottom section illustrates how the inference overlapping algorithm generates and blends the new patches, constructing $I^{\prime}$. It’s important to note that the image of the entire Canada region is $2540\times 1280$, where both dimensions are not divisible by $128$. As a result, the far right and bottom edges of the image contain regions that weren’t filled. For these smaller regions, we generate additional patches and apply the Inference Overlapping Algorithm to fill in the remaining pixels.
5 Conclusions
In this study, we presented a downscaling methodology based on a conditional Wasserstein GAN (WGAN-GP), leveraging low-resolution data from the GDPS and high-resolution data from the HRDPS, conditioned on high-resolution static covariates. We also conducted experiments incorporating the Frequency Separation approach to enhance the performance of the model. To implement the conditional WGAN-GP, we employed a UNET model as the generator. This UNET was designed to process two types of inputs: high-resolution static covariates and low-resolution GDPS data. Due to memory constraints, the training protocol relied on extracting random crops from single pairs of GDPS and HRDPS data. We carried out experiments to analyze the impact of including high-resolution static covariates in the model. Our findings indicate that these covariates significantly improved model convergence. Furthermore, experimental results demonstrated that the use of Frequency Separation contributed to further performance gains. We evaluated the generalization capability of the proposed model using a held-out test set. The results revealed that the best model configuration incorporated Frequency Separation, achieving favorable performance in terms of RMSE and Log Spectral Distance (LSD) metrics. To operationalize the model, we proposed a downscaling procedure for the entire Canadian domain. This procedure generates overlapping downscaled patches of size $128\times 128$ and blends them into the larger Canadian region using a linear interpolation approach. As future work, we aim to operationalize the model as a product capable of extending high-resolution wind forecast services from the current 48-hour horizon to the 10-day forecast horizon of the Canadian global forecast model. Additionally, we plan to investigate the application of recent state-of-the-art weather foundation models [37] for wind downscaling.

### v3

@@
 Abstract
-Wind downscaling is essential for improving the spatial resolution of weather forecasts, particularly in operational Numerical Weather Prediction (NWP).
+Downscaling based on Artificial Intelligence (AI) has been demonstrated to be a powerful tool for improving the spatial resolution of weather forecasts.
 This study advances wind downscaling by extending the DownGAN framework introduced by Annau et al.
-to operational datasets from the Global Deterministic Prediction System (GDPS) and High-Resolution Deterministic Prediction System (HRDPS), covering the entire Canadian domain.
-We enhance the model by incorporating high-resolution static covariates, such as HRDPS-derived topography, into a Conditional Wasserstein Generative Adversarial Network with Gradient Penalty, implemented using a UNET-based generator.
+to operational Canadian Numerical Weather Predictions (NWP), namely from the Global Deterministic Prediction System (GDPS) to the domain of the High-Resolution Deterministic Prediction System (HRDPS).
+We enhance the AI model by incorporating high-resolution static covariates, such as topography on the HRDPS grid, into a Conditional Wasserstein Generative Adversarial Network with Gradient Penalty, implemented using a UNET-based generator.
 Following the DownGAN framework, our methodology integrates low-resolution GDPS forecasts (15 km, 10-day horizon) and high-resolution HRDPS forecasts (2.5 km, 48-hour horizon) with Frequency Separation techniques adapted from computer vision.
-Through robust training and inference over the Canadian region, we demonstrate the operational scalability of our approach, achieving significant improvements in wind downscaling accuracy.
-Statistical validation highlights reductions in root mean square error (RMSE) and log spectral distance (LSD) metrics compared to the original DownGAN.
+Through robust training and inference over the whole Canadian region, we demonstrate the operational scalability of our approach, while achieving significant improvements in wind downscaling accuracy.
+Statistical validation highlights reduction in root mean square error (RMSE) and log spectral distance (LSD) metrics compared to the original DownGAN.
 High-resolution conditioning covariates and Frequency Separation strategies prove instrumental in enhancing model performance.
@@
 Accurate high-resolution wind forecasts are crucial for several application sectors such as aviation [1], wind energy [2], power infrastructure [3], public safety [4], wildfire management [5], avalanche warnings [6], winter sports [7], air quality [8] and water quality management [9].
-Typically, wind forecasts are obtained from numerical weather predictions (NWPs), starting with a global forecast model followed by a limited-area forecast model which takes the lower-resolution model as boundary conditions in a procedure called dynamical downscaling.
+Typically, high-resolution wind forecasts are obtained from numerical weather predictions (NWPs), starting with a global forecast model followed by a limited-area forecast model which takes the lower-resolution model as boundary conditions in a procedure called dynamical downscaling.
 In Canada, high-resolution numerical prediction models have a limited forecast horizon of 48 hours due to computational constraints.
@@
 [^1] Confusingly, what is called upsampling (downsampling) in the field of image processing is called downscaling (upscaling) in meteorology.
-Earlier methods were achieving an enhancement factor of 2 to 4 times, but with the advent of deep learning, an enhancement factor of 8 and more is now possible.
+Earlier image super-resolution methods were achieving an enhancement factor of 2 to 4 times, but with the advent of deep learning, an enhancement factor of 8 and more is now possible.
 Generative AI in particular has proven to be a powerful tool not only to artificially produce realistic-looking images and videos from a random seed, but also to condition the generative AI on coarser or partially missing data (called inpainting).
 AI-based techniques for downscaling weather and climate fields can be classified in an increasing order of complexity and recency.
-At the end of spectrum, ensemble decision trees techniques such random forests and extreme gradient boosting are the most similar to MLR and the easiest to implement.
+At the end of spectrum, ensemble decision trees [17] techniques such random forests and extreme gradient boosting are the most similar to MLR and the easiest to implement.
 Other classical machine learning techniques such as support vector machine and shallow artificial neural networks with a single hidden layer can also fit in that category.
-Methods based on Convolutional Neural Networks (CNNs) have proved to be both popular and powerful.
-Generative Adversarial Networks (GANs) are the first type of generative AI allowing to sample an ensemble of possible realizations [15].
-More recently, latent diffusion techniques have emerged as an interesting alternative to GAN.
+Methods based on Convolutional Neural Networks (CNNs) [18, 25, 26], have proved to be both popular and powerful.
+Generative Adversarial Networks (GANs) are the first type of generative AI allowing to sample an ensemble of possible realizations [15, 24, 27].
+More recently, latent diffusion techniques have emerged as an interesting alternative to GAN [28].
 Finally, methods based on vision transformers and in particular those that consist of fine-tuning a weather foundation model are emerging.
 In the literature, these methods have been applied to a variety of weather fields such as temperature, wind speed or wind vectors, and precipitation.
-Many of the published work focus on one or several small areas as a proof-of-concept, and few are used in an operational setting.
+Many of the published works focus on one or several small areas as a proof-of-concept, and few are used in an operational setting.
 It is important to note that similar AI techniques can be used either for downscaling weather forecasts, climate projections or reanalysis data, with the main difference in the choice and resolution of source and target data.
@@
 However, Frequency Separation methods adapted from computer vision and used by DownGAN show strong performance.
-The validation highlights the importance of high-resolution conditioning covariates and the utility of Frequency Separation in this context.
+The validation highlights the importance of high-resolution conditioning covariates and the utility of Frequency Separation for downscaling.
 The paper is divided as follows.
@@
 2 Dataset
-2.1 Global Deterministic Prediction System
+2.1 Numerical Weather Predictions
+Recall that the goal of this study is to extend the forecast horizon of the High-Resolution Deterministic Prediction System (HRDPS) by downscaling the Global Deterministic Prediction System (GDPS).
 The Global Deterministic Prediction System (GDPS) [29] is ECCC’s operational long-range forecast model.
 It represents the global configuration of the Global Environmental Multiscale (GEM) model [30] and it is run twice a day (at 00 UTC and 12 UTC), producing 10-day forecasts at a 15-km nominal resolution.
-2.2 High-Resolution Deterministic Prediction System
 The High-Resolution Deterministic Prediction System (HRDPS) [31] is a limited-area model (LAM) configuration of GEM for Canada and Northern United States.
@@
 The HRDPS is the main source of NWP guidance for the Meteorological Service of Canada for day 1 and day 2 forecasts.
-Figure 1: GDPS and HRDPS
 2.2 Predictors and Predictands
@@
 Low-resolution zonal and meridional surface wind from GDPS are taken as predictors along with surface temperature, as well as wind, temperature and vertical motion at 546 hPa.
-In contrast with [15], high-resolution geophysical fields are added as predictors.
+While we desired to stay close to the choices of predictors of [15], we did not include the Convective Available Energy Potential (CAPE) as it was not readily available in GDPS with an hourly resolution.
+We instead chose surface and mid-troposphere variables as they are expected to be important drivers of surface winds.
+However, we did not perform any variable selection experiment.
+In contrast with [15], the following three high-resolution geophysical fields are added as predictors: model orography, land-water mask and surface roughness length.
 These static high-resolution covariates can be used as predictors since these variables are constant at the time-scale of weather prediction.
-The topography field is particularly important in mountainous regions where wind channeling is predominant.
+The model orography field is particularly important in mountainous regions where wind channeling is predominant.
 The land-water mask not only provides an indication of changing surface roughness between the smoother water and other land, but can also be a driver of local winds such as sea breezes and lake breezes.
-Finally, the surface roughness variable modulates the speed of the wind and changing surface roughness can drive turbulence in the planet boundary layer.
+Finally, the surface roughness variable modulates the speed of the wind, and changes in this covariate can drive turbulence in the planetary boundary layer.
 Table LABEL:table:vars_description describes the meteorological variables used for training the AI downscaling model.
-Table 1: Description of Variables for GDPS and HRDPS
+Table 1: Description of predictors and predictands for the downscaling experiments.
+The HRDPS static covariates are extra high-resolution predictors used in the Conditional WGAN-GP, whereas the DownGAN baseline we implemented only uses GDPS predictors.
 Description
 GDPS Predictors
-True geographical West-East (zonal) component of the horizontal wind at the surface (10 m) [m/s].
-True geographical South-North (meridional) component of the horizontal wind at the surface (10 m) [m/s].
-Air temperature at the surface (1.5 m) [°C].
-Air temperature vertically interpolated at 546 hPa [°C].
-True geographical West-East (zonal) component of the horizontal wind vertically interpolated at 546 hPa [m/s].
-True geographical South-North (meridional) component of the horizontal wind vertically interpolated at 546 hPa [m/s].
-Vertical motion vertically interpolated at 546 hPa [Pa/s].
+$\mbox{U}_{\mbox{surf}}$ True geographical West-East (zonal) component of the horizontal wind at the surface (10 m) [m/s].
+$\mbox{V}_{\mbox{surf}}$ True geographical South-North (meridional) component of the horizontal wind at the surface (10 m) [m/s].
+$\mbox{T}_{\mbox{surf}}$ Air temperature at the surface (1.5 m) [°C].
+$\mbox{T}_{546}$ Air temperature vertically interpolated at 546 hPa [°C].
+$\mbox{U}_{546}$ True geographical West-East (zonal) component of the horizontal wind vertically interpolated at 546 hPa [m/s].
+$\mbox{V}_{546}$ True geographical South-North (meridional) component of the horizontal wind vertically interpolated at 546 hPa [m/s].
+$\mbox{W}_{546}$ Vertical motion vertically interpolated at 546 hPa [Pa/s].
 HRDPS Predictands
-True geographical West-East (zonal) component of the horizontal wind at the surface (10 m) [m/s]. Denoted by u10.
-True geographical South-North (meridional) component of the horizontal wind at the surface (10 m) [m/s]. Denoted by v10.
+u10 True geographical West-East (zonal) component of the horizontal wind at the surface (10 m) [m/s].
+v10 True geographical South-North (meridional) component of the horizontal wind at the surface (10 m) [m/s].
 HRDPS Static Covariates
-Model orography [m].
-Water/land mask [fraction].
-Roughness length [m].
+me Model orography [m].
+mg Water/land mask [fraction].
+z0 Roughness length [m].
 2.3 Data pre-processing
 The map projection of the GDPS outputs is a Yin-Yang grid with a different rotated latitude-longitude map projection than the one of the HRDPS grid.
-We regrid the GDPS grid to the HRDPS grid using nearest neighbor interpolation.
+To align the grids, we project [^2] The regridding process was done using the xESMF Python Library [32].
+the GDPS grid to the HRDPS grid using nearest neighbor interpolation.
 The interpolated data is then reduced by a factor of 8 (20-km nominal resolution) as the input of the AI downscaling method.
@@
 1 illustrates the respective domains of GDPS (cropped over North-America) and of HRDPS.
+Figure 1: Example of zonal and meridional winds on the GDPS and HRDPS grids. Note that the GDPS grid was cropped over North America. The domain of HRDPS grid is indicated by a black rectangle.
 3 Methods
-In this study, we adopt the deep learning framework for downscaling wind components proposed by [15].
+In this study, we adapt the deep learning framework for downscaling wind components proposed by [15].
 This framework is based on Wasserstein Generative Adversarial Networks with Gradient Penalty (WGAN-GP) [33] and incorporates elements from Generative Adversarial Networks for image super-resolution (SRGAN) [34] as a baseline for the downscaling task.
-In the following sections, we elaborate on the application of WGANs for the downscaling problem, describe the model architecture, and introduce our approach to integrating high-resolution static covariates into the training process using UNETs [31].
+In the following sections, we elaborate on the application of WGANs for the downscaling problem, describe the model architecture, and introduce our approach for integrating high-resolution static covariates into the training process using UNETs [35].
 3.1 Conditional WGAN-GP for Downscaling
-Given a training set consisting of dynamic low-resolution (GDPS) and high-resolution (HRDPS) forecast hours, denoted as ${\bf x}$ and ${\bf y}$, respectively.
-A conditional WGAN-GP, conditioned on the covariates ${\bf y_{cov}}$ in the HRDPS domain, aims to minimize the following losses for the critic $\mathcal{L}_{C}$ and generator $\mathcal{L}_{G}$ [15, 29, 32]:
+We consider a training set consisting of dynamic low-resolution predictors, such as GDPS fields, and high-resolution predictands, such as HRDPS fields, with matched forecast hours, which we denote as ${\bf x}$ and ${\bf y}$, respectively.
+A conditional WGAN-GP, conditioned on the covariates ${\bf y_{cov}}$ in the HRDPS domain, aims to minimize the following losses for, respectively, the critic $\mathcal{L}_{C}$ and generator $\mathcal{L}_{G}$ [15, 33, 36]:
 $\mathcal{L}_{C}=\mathbb{E}_{G({\bf x})\sim\mathbb{P}_{g}}[C(G({\bf x}\mid{\bf y_{cov}}))]-\mathbb{E}_{{\bf y}\sim\mathbb{P}_{r}}[C({\bf y})]+\lambda\mathbb{E}_{{\bf\hat{y}}\sim\mathbb{P}_{\hat{y}}}[(\|\nabla_{\hat{y}}C(\hat{y})\|_{2}-1)^{2}]$ (1)
-$\mathcal{L}_{G}=-\gamma\mathbb{E}_{G({\bf x})\sim\mathbb{P}_{g}}[C(G({\bf x}\mid{\bf y_{cov}}))]+\alpha\mathbb{E}_{{\bf y}\sim\mathbb{P}_{r}}l_{c}({\bf y},G({\bf x}\mid{\bf y_{cov}}))$ (2)
-In this formulation, $C$ denotes a real-valued function known as the Critic, which is defined on the HRDPS domain.
-The function $G$ represents the Generator, which maps low-resolution GDPS to the HRDPS domain.
-The symbol $\mathbb{P}_{g}$ denotes the conditional distribution of the generated downscaled GDPS data given the covariates ${\bf y_{cov}}$, while $\mathbb{P}_{r}$ represents the distribution of the HRDPS points.
+and
+$\mathcal{L}_{G}=-\gamma\mathbb{E}_{G({\bf x})\sim\mathbb{P}_{g}}[C(G({\bf x}\mid{\bf y_{cov}}))]+\alpha\mathbb{E}_{{\bf y}\sim\mathbb{P}_{r}}l_{c}({\bf y},G({\bf x}\mid{\bf y_{cov}})).$ (2)
+In this formulation, $C$ denotes a real-valued function known as the Critic, which is defined on the high-resolution domain.
+The function $G$ represents the Generator, which maps low-resolution predictors to the high-resolution predictands.
+The symbol $\mathbb{P}_{g}$ denotes the conditional distribution of the generated downscaled data given the high-resolution static covariates ${\bf y_{cov}}$, while $\mathbb{P}_{r}$ represents the distribution of the high-resolution predictands.
 The expression $\mathbb{E}_{{\bf\hat{y}}\sim\mathbb{P}_{\hat{y}}}[(\|\nabla_{\hat{y}}C(\hat{y})\|_{2}-1)^{2}]$ represents the gradient penalty, which is incorporated into the formulation to ensure valid Wasserstein distances by enforcing 1-Lipschitz continuity for the Critic.
@@
 Here, $B$ represents the batch size, $W$ and $H$ denote the spatial width and height, respectively, while $C$ indicates the number of input channels (variates).
-The subscripts $l$ and $h$ refer to the GDPS and HRDPS datasets, respectively.
+The subscripts $l$ and $h$ refer to the high- and low-resolution datasets, respectively.
 Additionally, ${\bf y_{cov}}$ is a tensor with dimensions $C_{cov}\times W_{l}\times H_{l}$, where values are replicated across the batch size $B$.
@@
 For a more comprehensive understanding of these concepts, please refer to the relevant literature [15, 33, 37].
-3.1.1 Frequency Separation
+3.2 Frequency Separation
 The core idea of the frequency separation approach is to decompose the low- and high-frequency components and treat them differently during WGAN-GP training [15, 38], with the aim of enhancing the quality of downscaled wind components.
 This allows the adversarial training to focus specifically on generating the high-frequency components, which are crucial for capturing fine details.
-In this approach, the high-frequency components are derived as ${\bf y}_{h}={\bf y}-H({\bf y})$ for the HRDPS data, and $G({\bf x}\mid{\bf y_{cov}})_{h}=G({\bf x}\mid{\bf y_{cov}})-H(G({\bf x}\mid{\bf y_{cov}}))$ for the generated downscaled data, where $H$ represents a low-pass filter.
-In this case, during WGAN-GP training, the Critic will focus on optimizing the loss based only in the high frequency components ${\bf y}_{h}$ and $(G({\bf x}\mid{\bf y_{cov}})_{h}$ instead of ${\bf y}$ and $(G({\bf x}\mid{\bf y_{cov}}))$.
-On the other side the Generator delegates the generation of high frequency components to the adversarial component $-\mathbb{E}_{G({\bf x})_{h}\sim\mathbb{P}_{g}}[C(G({\bf x}\mid{\bf y_{cov}}))_{h}]$ and the low frequency components to the content loss $\mathbb{E}_{{\bf y}\sim\mathbb{P}_{r}}l_{c}(H({\bf y}),H(G({\bf x}\mid{\bf y_{cov}})))$.
-In [15] is proposed the partial frequency separation, this approach is similar than before but the adversarial loss is used in its original formulation, i.e., $-\mathbb{E}_{G({\bf x})\sim\mathbb{P}_{g}}[C(G({\bf x}\mid{\bf y_{cov}}))]$
+In this approach, the high-frequency components are derived as ${\bf y}_{high}={\bf y}-\mathscr{L}({\bf y})$ for the HRDPS data, and $G({\bf x}\mid{\bf y_{cov}})_{high}=G({\bf x}\mid{\bf y_{cov}})-\mathscr{L}(G({\bf x}\mid{\bf y_{cov}}))$ for the generated downscaled data, where $\mathscr{L}$ represents a low-pass filter.
+In this case, during WGAN-GP training, the Critic will focus on optimizing the loss based only in the high frequency components ${\bf y}_{high}$ and $G({\bf x}\mid{\bf y_{cov}})_{high}$ instead of ${\bf y}$ and $G({\bf x}\mid{\bf y_{cov}})$.
+On the other side the Generator delegates the generation of high frequency components to the adversarial loss component $-\mathbb{E}_{G({\bf x})_{high}\sim\mathbb{P}_{g}}[C(G({\bf x}\mid{\bf y_{cov}})_{high})]$ and the low frequency components to the content loss $\mathbb{E}_{{\bf y}\sim\mathbb{P}_{r}}l_{c}(\mathscr{L}({\bf y}),\mathscr{L}(G({\bf x}\mid{\bf y_{cov}})))$.
+[^3] Annau et.al., proposed partial frequency separation [15], this approach is similar to frequency separation, but the adversarial loss is used in its original formulation, i.e., $-\mathbb{E}_{G({\bf x})\sim\mathbb{P}_{g}}[C(G({\bf x}\mid{\bf y_{cov}}))]$.
 3.3 Architectures
@@
 The labels below the blocks denote the kernel size $k$, the number of filters $n$, and the stride $s$ used in each layer.
-The Critic utilizes a VGG-style architecture that progressively reduces the spatial dimensions of the input, which may include HRDPS wind components or the downscaled version of GDPS wind components provided by the Generator network, through a sequence of convolutional layers.
-The network processes the input data using 3x3 convolutions with LeakyReLU activations, where each layer increases the number of feature maps while downsampling the spatial dimensions using a stride of 2.
+The Critic utilizes a VGG-style architecture that progressively reduces the spatial dimensions of the input [39], which may include HRDPS wind components or the downscaled version of GDPS wind components provided by the Generator network, through a sequence of convolutional layers.
+The network processes the input data using 3×3 convolutions with LeakyReLU activations [^4] LeakyReLU are defined as: $\sigma(x)=\max(x,\alpha x)$ where $x$ is the input to the activation function, and $\alpha$ is a small positive parameter that determines the slope for negative inputs., where each layer increases the number of feature maps while downsampling the spatial dimensions using a stride of 2.
 This approach allows the network to expand the number of channels up to eight times the input dimension.
 The extracted features are then passed through a final processing block, a regression-style scoring mechanism that reduces the feature vector to a single scalar output.
-This scalar score, used as an approximation of the Wasserstein distance, indicates the Critic’s assessment of how real the input data appears.
-Figure 2: Schematic of the Critic Network
+This scalar score, used as an approximation of the Wasserstein distance, indicates the Critic’s assessment of the physical realism of the generated data.
+Figure 2: Schema of the Critic Network.
+The input consists of either a block of high-resolution NWP predictands or downscaled data generated by the Generator.
+The output is a real-valued score representing the Critic’s confidence that the input originates from actual NWP data—a higher score indicates a more physically realistic appearance.
+The labels $k$, $n$ and $s$ below the blocks denote the kernel size, the number of filters, and the stride used in each layer, respectively.
+The numbers above the blocks represent the tensor dimensions as a function of the number of feature maps, width, and height.
 3.3.2 Generator Network
 The generator network implemented in the proposed model draws upon concepts from [15, 34] and is structured according to a U-Net architecture.
-The architecture is depicted in Figure 3 We employ this design because it facilitates the integration of HRDPS covariates and GDPS variates as inputs.
+The architecture is depicted in Figure 3.
+We employ this design because it facilitates the integration of static HRDPS covariates and GDPS variates as inputs.
 The Generator embeds the HRDPS covariates through a sequence of downsampling layers reducing the spatial dimensions by a factor of eight.
@@
 This combined tensor passes through a series of residual-in-residual dense blocks, for a deep feature extraction.
-Following this stage, the network employs a sequence of upsampling layers that utilize pixel shuffle operations, enabling the progressive restoration of spatial resolution.
+Next, the network applies a series of upsampling layers that leverage pixel shuffle operations [40] to progressively restore spatial resolution.
+The pixel shuffle operation restructures the input tensor’s feature maps into a higher-resolution grid, effectively increasing spatial resolution.
+For an upscale factor of 2, the operation requires four times the number of input feature maps to redistribute them into the upsampled spatial dimensions while maintaining the number of output channels.
+Figure 3 shows an $n=64$ ($n64$ below each upsampling block), representing 64 input feature maps required to produce an upsampled output tensor with 16 feature maps.
 Throughout the upsampling process, skip connections from the downsampling layers are leveraged to facilitate the merging of local and global features while preserving high-frequency details.
-The final output layer incorporates a convolutional operation to generate the GDPS downscaled version of wind componets.
-This U-Net architecture serves as the mechanism for implementing the conditional Wasserstein GAN, effectively utilizing HRDPS covariates to guide the generative process GDPS wind components.
-Figure 3: Schematic of the Generator Network
+The final output layer incorporates a convolutional operation to generate the GDPS downscaled version of wind components.
+This U-Net architecture serves as the mechanism for implementing the conditional Wasserstein GAN, effectively utilizing static HRDPS covariates to guide the generative process of downscaled GDPS wind components.
+Figure 3: Schematic of the Generator Network.
+The inputs consist of static high-resolution covariates (top-left) and low-resolution predictors (buttom-left).
+The static covariates go through three successive coarsening passes and are then concatenated with the low-resolution predictors.
+The resulting combination then goes sixteen times through the core component called ”residual in residual dense block”.
+Three refinement passes with skip connections followed by a last group of convolution yield the final high-resolution output.
 4 Experiments
 4.1 Dataset Split
-This section outlines the experimental setup, using a dataset of matched GDPS and HRDPS forecast hours.For each prediction system, 24 hours of forecasts are used for each day consisting of forecast lead-times 6-17 hours for each of the 00 and 12 UTC runs.
+This section describes the experimental setup, using a dataset of matched GDPS and HRDPS forecast hours.
+For each prediction system, 24 hours of forecasts are used for each day consisting of forecast lead-times 6-17 hours for each of the 00 and 12 UTC runs.
 In this way, we avoid both the model spin-up time, and later forecast times where the two systems might have diverged because of error growth.
 For training, we selected all GDPS and HRDPS files with predictions starting between July 2022 to June 2023.
-The dataset was randomly split into 7,150 files for training, 1,178 for validation, and 1,162 for testing.
+The dataset was randomly split into 7 150 files for training and 1 178 for validation.
+We use an independent dataset, separate from the training and validation intervals, as the test set.
+It comprises forecast lead times from 6 to 17 hours for July 2023 (682 files), August 2023 (682 files), January 2024 (618 files), and February 2024 (696 files).
 4.2 Metrics
@@
 Additionally, we employed the Radially Averaged Log-Spectrum Distance (LSD) [19]:
-$LSD(\overline{P}_{true},\overline{P}_{pred})=\sqrt{\frac{1}{N_{r}}\sum_{r}\left(10\log_{10}\frac{\overline{P}_{true}(r)}{\overline{P}_{pred}(r)}\right)^{2}},$ (3)
-where $P(r)$ is the Radially Averaged Power Spectrum Density [19] Observe that LSD quantifies the difference between predicted and true power spectra on a logarithmic scale.
-Subscripts pred and true are the $P(r)$ of predicted and true wind fields, respectively.
+$LSD(\overline{P}_{ref},\overline{P}_{pred})=\sqrt{\frac{1}{N_{r}}\sum_{r}\left(10\log_{10}\frac{\overline{P}_{ref}(r)}{\overline{P}_{pred}(r)}\right)^{2}},$ (3)
+where $P(r)$ is the Radially Averaged Power Spectrum Density (RAPSD) [19, 41] Observe that LSD quantifies the difference between predicted and true power spectra on a logarithmic scale.
+Subscripts pred and ref are the $P(r)$ of predicted and reference wind fields, respectively.
 4.3 Training Procedures
-The GDPS and HRDPS datasets cover the entire Canadian region.
-We applied a nearest-neighbor regridding approach to align the GDPS data with the HRDPS grid of $(2540,1280)$ spatial values, because each data set was provided in different rotated coordinate systems [^2] The regridding process was done using the xESMF Python Library [35].
-We use this regridding because preserves original pixel values avoiding interpolation artifacts.
-In this sense, each GDPS and HRDPS file is represented by tensors of sizes$(7,2540,1280)$ and $(2,2540,1280)$, where 7 and 2 represent the variable counts in GDPS and HRDPS, respectively, as detailed in Table LABEL:table:vars_description.
+The regridded GDPS (see Section 2.3) and HRDPS datasets cover the entire Canadian region (except the Arctic Archipelago) with a grid of $(2540,1280)$ spatial values.
+In this sense, each GDPS and HRDPS file is represented by tensors of sizes $(7,2540,1280)$ and $(2,2540,1280)$, where 7 and 2 represent the variable count in GDPS and HRDPS, respectively, as detailed in Table LABEL:table:vars_description.
 To train the WGAN-GP, we based our implementation in the open-source implementation from [42].
@@
 We used spatial random cropping for training with a $(128\times 128)$ crop size and a downsampling factor of 8 for the regridded GDPS, resulting in GDPS tensors of $(7,16,16)$ for each corresponding HRDPS tensor of $(2,128,128)$.
-We generated 832 crops per GDPS/HRDPS pair, amounting to 5,948,800 random crops for the 7,150 pairs in the training dataset.
-Here, one ”epoch” corresponds to using all 832 random crops from a single GDPS-HRDPS forecast hour.
-This random crop setup allows six generator updates using a batch size of 32.
+We generated 832 crops per GDPS/HRDPS pair, amounting to 5 948 800 random crops for the 7 150 pairs in the training dataset.
+Here, one ”training step” corresponds to using all 832 random crops from a single GDPS-HRDPS forecast hour.
+This random crop setup allows six generator updates using a batch size of 32, fitting in the 40 GB of RAM of a Nvidia’s A100 GPU chip we used for training.
 The UNET generator network is configured with inputs of $(32,7,16,16)$ (batch size, variables, spatial dimensions) and receives an HRDPS static covariate tensor of $(3,128,128)$.
@@
 The critic network processes input tensors of $(32,2,128,128)$, derived from either HRDPS data or the downscaled GDPS output.
-The proposed conditional WGAN-GP was trained for 65,250 epochs, with validation involving 32 random crops per GDPS-HRDPS file in the validation set.
+The proposed conditional WGAN-GP was trained for 65,250 training steps, with validation involving 32 random crops per GDPS-HRDPS pair in the validation set.
 4.4 Effect of using HRDPS Static Covariates
 In this experiment, we investigate the impact of incorporating HRDPS static covariates on downscaling accuracy.
-To this end, we compare the DownGAN model [15] using the open-source implementation from [36], against a the proposed conditional WGAN-GP that utilizes HRDPS covariates.
+To this end, we compare the DownGAN model [15] using the open-source implementation from [42], against the proposed conditional WGAN-GP that utilizes HRDPS static covariates.
 Both models are trained with identical hyperparameters and configurations to ensure a fair comparison.
 The primary difference lies in their generator architectures: DownGAN employs a generator based on the SRGAN network introduced by [34], while the WGAN-GP model adapts the DownGAN generator to a UNET structure, allowing it to effectively incorporate the additional HRDPS covariates.
-Figure 4: Training and validation curves: DownGAN vs conditional WGAN-GP
-Figure 4 presents the MSE curve trajectories for the DownGAN and the Conditional WGAN-GP models across training and validation epochs.
-The MSE values are averaged over intervals of 7250 epochs, which correspond to the number of forecast hours in the training set.
+Figure 4: Training and validation curves: DownGAN vs conditional WGAN-GP. Conditional WGAN-GP outperforms DownGAN in terms of the validation MSE curve and achieves faster convergence during training.
+Figure 4 presents the MSE curve trajectories for the DownGAN and the Conditional WGAN-GP models across training and validation steps.
+The MSE values are averaged over intervals of 7 250 training steps, which correspond to the number of files in the training set.
 In the plot, the DownGAN model’s training and validation MSE average values are represented by blue and red markers, respectively, while the Conditional WGAN-GP model’s MSE average values are shown in green (training) and orange (validation).
@@
 The minimum validation MSE average value for each model is marked with an ”x”.
-The Conditional WGAN-GP model attains a lower validation error across all intervals, suggesting its enhanced capability for downscaling accuracy.
+The Conditional WGAN-GP model attains a lower validation error across all intervals, suggesting its enhanced capability for accurately downscaling surface winds.
 4.5 Effect of Frequency Separation
@@
 Additionally, we included a conditional WGAN-GP model without FS or PFS as a baseline (i.e., same MSE validation curve from Figure 4).
-For the experiments, we applied transfer learning starting from epoch 21,750, and continuing training up to epoch 50,750.
+For the experiments, we applied transfer learning starting from training step 21 750, and continuing training up to step 50 750.
 Figure 5 presents a comparison of the models’ performance based on the MSE validation metric.
-Figure 5: Effect of Frequency separation in the validation curve
+Figure 5: Effect of Frequency separation in the validation curve.
+Frequency separation with filter sizes of (5,5), (9,9) or (13,13) has a positive effect on the MSE validation score (0.86, 0.87 and 0.9, respectively), but partial frequency separation does not improve the scores or only slightly (MSE validation score of 0.93) for filter size of (5,5).
 From the experiments, we observed that FS is in general better than PFS, and than small filter kernel sizes are better in term of the validation MSE metric.
 For this experiment, WGAN-GP with FS achieved the best configuration using a filter size of $(5,5)$, obtaining an average MSE validation score of 0.86 in the interval $[43500,50750]$.
-This result outperformed both the best PFS configuration (filter size of $(5,5)$ and MSE validation score of 0.93) and the case without frequency separation (MSE validation score of 0.97)
-Figure 6: Effect of Frequency separation in downscaling at epoch 36250 — Left to right images: GDPS, GDPS downscaled with no frequency separation case, GDPS downscaled with WGAN-GP with FS using filter $(7,7)$ and HRDPS reference
-Figure 6 shows the effect of FS in dealing with high frequencies as a function of epochs 29000, 36250, 43500 and 50750.
-In this case, checkboard artifacts are less evident in WGAN-GP with FS than in the case with no frequency separation.
+This result outperformed both the best PFS configuration (filter size of $(5,5)$ and MSE validation score of 0.93) and the case without frequency separation (MSE validation score of 0.97).
+Figure 6: Effect of frequency separation on downscaling the v10 wind component at 15:00 UTC on July 1, 2023, at training steps 29 000, 36 500, 43 500, and 50 750.
+Rows show different configurations; columns show training steps.
+Corresponding wind fields from GDPS and HRDPS are shown for reference.
+All the downscaled fields capture well the main features of the HRDPS forecast, but frequency separation reduces checkboard artifacts and a larger filter size leads to a sharper forecast.
+Figure 6 shows the effect of FS in dealing with high frequencies as a function of training steps 29 000, 36 250, 43 500 and 50 750.
+In this case, checkboard artifacts are less evident in WGAN-GP with FS than in the case with no frequency separation (NFS), which can be explained by the greater focus on high-frequency components when training with FS.
 4.6 Metrics on the Test Set
-This section presents the metrics computed on a test set comprising 1,162 GDPS forecast hours.
-Figure 7 displays the results, contrasting the 1,162 downscaled GDPS forecast hours against HRDPS references in terms of RMSE and LSD metrics.
-We used Bilinear interpolation and nearest neighbor (NN) interpolation as baseline methods.
-Additionally, we evaluated several models, including the DownGAN model at checkpoint 29000, the Conditional WGAN-GP model at checkpoints 21,750 and 43,500 (no frequency separation scenario: NFS - 21,750 and NFS - 43,500, respectively), and the Conditional WGAN-GP model with frequency separation with kernels $(13,13),(9,9)$ and $(5,5)$ at checkpoints 50,750 denoted by FS13 - 50,750, FS9 - 50,750, and FS5 - 50,750 respectively.
-The decision to select a specific model checkpoint was based on the lowest validation MSE metric (See experiments from Section 4.5).
-Figure 7: RMSE and LSD Metrics on the Test Set
-As shown by Figure 7, based on the test error in both metrics for the two wind components, the interpolation methods performed the worst.
-Additionally, the DownGAN model exhibited worse performance than the Conditional WGAN-GP variations.
-Among the Conditional WGAN-GP models, the method with frequency separation and a kernel size of $5,5$ provided the best performance in terms of the RMSE metric, and the one with kernel size of $9,9$ in terms of the LSD metric Table LABEL:tab:metrics_summary presents a summary of the metrics on the test set in terms of their mean and standard deviation statistics.
+This section presents the metrics computed on the independent test set described in Section 2.
+We compute the metrics by performing downscaling over the entire Canadian domain using the trained generator.
+Although the models were trained with random 16 × 16 crops to produce images of 128 x 128, we use the trained generator to perform downscaling on the entire Canadian region.
+Specifically, we send the full GDPS file, which covers the entire Canadian domain, directly to the generator without the need to manually stitch together the 128 x 128 blocks [^5] This is possible because a convolutional layer’s parameter count depends only on the filter size and the number of input and output feature maps, remaining unaffected by input width and height.
+This enables convolutional networks to handle varying spatial dimensions without changing the number of trainable parameters..
+To ensure compatibility with the network’s design for an exact 8x resolution increase, we adjust the original dimensions.
+The original GDPS input dimensions (162, 318) are modified to (161, 317), producing a downscaled GDPS output with dimensions of (1288, 2536).
+Similarly, we adjust the original HRDPS reference dimensions from (1290, 2540) to (1288, 2536) to enable direct comparison with the downscaled GDPS version.
+Figure 7 presents an example of downscaling the u10 wind component from the GDPS hourly forecast from the test set, corresponding to the 12:00 UTC run at 06:00 on February 22, 2024, over the entire Canadian domain, using the WGAN-GP model with a filter size of (13,13) and checkpoint 50 750.
+In this case, the RMSE between the downscaled GDPS and HRDPS is 0.842, while the log spectral distance (LSD) is 0.655.
+Subjective assessment of the downscaled GDPS confirms that overall the output matches closely the HRDPS and enhances the details in mountainous areas and sharpen the boundary between weak and strong winds over lakes.
+However, some details are still smoothed or mismatched over the two oceans (Atlantic and Pacific).
+The use of static covariates thus enable to effectively downscale topographically-driven features, but in the absence of strong topographic features such as over the oceans and in the prairies, the AI downscaling model only partially reconstructs high-frequency details.
+Over the mountains, the power spectra match very well, but over the prairies the power spectra of the downscaled GDPS loses some energy below 50 km resolution.
+Over the oceans, the detailed locations of the wind features are wrong even if the features themselves look fairly realistic.
+The power spectra thus match very well even if the wind patterns are different.
+For this particular winter case, the wind over lakes is stronger for the downscaled GDPS than for the target HRDPS, possibly because a failure of the downscaling model to account for a change of surface roughness length when lakes ice.
+It is reflected with a higher energy power spectra at higher frequencies for the downscaled GDPS compared to the target HRDPS.
+Figure 7: GDPS, HRDPS, and GDPS downscaled forecasts for the test set (u10 wind component) corresponding to the 12:00 UTC run at 06:00 on February 22, 2024.
+The downscaling was performed over the entire Canadian domain using the WGAN-GP model with frequency separation and a filter size of (13,13), utilizing trained model checkpoint 50 750.
+The RAPSD graph show the RAPSD plots for HRDPS the downscaled GDPS, and the GDPS, here the gray dashed vertical line describe the cut frequency for the GPDS $20km^{-1}$.
+Bounding boxes in black over GDPS are visualized in bottom part, for four distinct regions: Mountains, Prairies, Lakes, and Oceans.
+This figure highlights how the AI downscaling model captures fine-scale features in areas with strong topographic influence while revealing its limitations in regions with no topographic information, such as the oceans.
+Figures 8 and 9 shows the distribution of RMSE and LSD values on the test set, comparing downscaled GDPS forecasts against HRDPS references.
+The evaluation is conducted separately for the u10 and v10 components and further stratified by month.
+Bilinear interpolation and nearest-neighbor (NN) interpolation serve as baseline methods.
+In addition, we assess multiple models, including the DownGAN model at checkpoint 29 000 and the Conditional WGAN-GP model under different configurations: this model is evaluated at checkpoints 21 750 and 43 500 in the absence of frequency separation (denoted as NFS-21 750 and NFS-43 500, respectively) and with frequency separation using filter sizes of $(13,13)$, $(9,9)$, and $(5,5)$ at checkpoint 50 750 (denoted as FS13-50 750, FS9-50 750, and FS5-50 750, respectively).
+The selection of model checkpoints is guided by the lowest validation MSE observed in the experiments detailed in Section 4.5.
+Figure 8: Violin plots of the RMSE of the u10 and v10 on the test sets for four different months: July 2023, August 2023, January 2024 and February 2024.
+Simple interpolation such as bilinear and nearest neighbour performs worst for all months and wind components.
+The DownGAN model performs slightly better, but the Conditional WGAN-GP is best.
+Training for 43 500 steps is slightly better than for 21750 steps and adding frequency separation further enhances the scores.
+Overall, RMSE is lower for Winter months (January-February) than for Summer months (July-August) and similar for u10 and v10 wind components.
+Figure 9: Violon plots of LSD on the test sets for the same months as Fig.
+8.
+Bilinear interpolation has clearly the worst power spectrum for all months and wind components followed by the nearest neighbor interpolation.
+DownGAN does similar than the Conditional WGAN-GP without frequency separation, except for v10 in Summer months where it does better.
+However, the power spectra that are the most similar to the HRDPS power spectra are the Conditional WGAN-GP with frequency separation.
+Frequency separation with large filter size of (13,13) is better than with smaller filter size of (5,5), particularly for July and August.
+Analyzing the RMSE distributions in Figure 8 for both wind components, u10 and v10, we observed that the interpolation methods (Bilinear and NN) produced the highest RMSE values, indicating their inferior performance.
+The DownGAN model also yielded higher RMSE values compared to the Conditional WGAN-GP variations.
+Among the Conditional WGAN-GP models, those incorporating frequency separation achieved the lowest RMSE, demonstrating superior performance in this metric.
+The case of LSD distributions shown in Figure 9 highlights the role of the filter size in the Conditional WGAN-GP with Frequency Separation.
+In this case, a filter size of $13,13$ or $9,9$ achieved the lowest LSD metric, indicating its superior performance according to this metric.
+Table LABEL:tab:metrics_summary presents a summary of the metrics in the test set, aggregating all the months, in terms of their median and median absolute deviation (MAD) statistics.
+Marked in bold is the best model for RMSE and for LSD, for both u10 and v10.
+Overall, the Conditional WGAN-GP with frequency separation with a filter size of (5,5) has the lowest RMSE for both u10 and v10, but increasing the filter size to (9,9) or (13,13) leads to, respectively, the lowest LSD for v10 and u10.
+Note that the difference in median RMSE between these three models is only 0.04 m/s, which is smaller than the MAD of 0.09 m/s.
+However, the difference of LSD for u10 is 0.38 dB, which is more than the MAD of 0.18 dB for FS13.
+Similarly the difference of LSD for v10 is 0.16 dB which is about the same as the MAD of F13, but the difference between FS9 and FS13 is only 0.004 dB.
+The F13 model offers the best balance between RMSE and LSD performance.
 Table 2: Summary of results.
 Method RMSE (u10) RMSE (v10) LSD (u10) LSD (v10)
-Mean Std Mean Std Mean Std Mean Std
-DownGAN - 30000 0.953 0.333 0.904 0.305 2.287 1.223 2.327 1.292
-FS13 - 50750 0.785 0.296 0.785 0.293 1.820 0.939 1.852 0.980
-FS9 - 50750 0.771 0.298 0.764 0.286 1.755 0.914 1.671 0.887
-FS5 - 50750 0.749 0.296 0.744 0.287 1.842 0.964 1.756 0.934
-NFS - 43500 0.803 0.311 0.778 0.294 1.889 0.973 1.847 0.970
-NFS - 21750 0.843 0.326 0.812 0.303 1.962 0.992 2.082 1.078
-Bilinear 1.012 0.396 0.993 0.378 3.462 2.379 3.476 2.412
-KNN 1.041 0.399 1.023 0.383 3.994 0.790 4.008 0.764
-4.7 Inference over the whole Canadian Region
-Downscaling is crucial for analyzing large geographical regions; however, memory constraints can arise due to factors such as spatial resolution, data quantity, and model size.
-To address this issue, downscaling is often performed on smaller regions or patches, which are then stitched together to create a larger image.
-To downscale wind components on the entire Canadian domain we used the conditional WGAN-GP to generate overlapping $128\times 128$ images across the domain and then blend these overlapping areas using a weight matrix $W$, which is predefined $128\times 128$ matrix with weights in the interval $[0,1]$, increasing radially toward the center.
-Specifically, to generate a larger downscaling image $I$ from smaller $128\times 128$ images (patches), the algorithm starts by generating consecutive non-overlapping patches with a stride of 128 and copying the result directly into the corresponding region of $I$.
-Then, the entire region is again parsed, but this time the new patches overlap the artefacts.
-The final image is created by using linear interpolation between the patches generated in the last step and the correspondent pixels that were already in $I$ by using the weight matrix $W$.
-Each value for pixel $ij$ in image I’ is given by $I_{ij}^{\prime}=I_{ij}*(1-W_{ij})+\text{patch}_{ij}*W_{ij}$ Figure 8 provides a visual representation of this approach.
-Figure 8: Inference overlapping strategy used to downscaling on the whole Canadian Region
-The top part of the figure highlights the region $I$, which contains artifacts and is constructed using consecutive non-overlapping patches.
-The bottom section illustrates how the inference overlapping algorithm generates and blends the new patches, constructing $I^{\prime}$.
-It’s important to note that the image of the entire Canada region is $2540\times 1280$, where both dimensions are not divisible by $128$.
-As a result, the far right and bottom edges of the image contain regions that weren’t filled.
-For these smaller regions, we generate additional patches and apply the Inference Overlapping Algorithm to fill in the remaining pixels.
+Median MAD Median MAD Median MAD Median MAD
+Bilinear 1.224 0.103 1.220 0.107 4.703 1.418 4.124 1.083
+NN 1.252 0.103 1.252 0.110 2.625 0.417 2.843 0.398
+DownGAN-30000 1.142 0.080 1.116 0.100 1.743 0.566 1.408 0.334
+FS13-50750 0.993 0.086 1.023 0.100 0.744 0.181 0.735 0.164
+FS9-50750 0.974 0.087 0.995 0.109 0.877 0.253 0.731 0.188
+FS5-50750 0.959 0.089 0.981 0.111 1.124 0.315 0.998 0.220
+NFS-21750 1.048 0.097 1.051 0.106 1.689 0.539 1.985 0.498
+NFS-43500 0.997 0.090 1.015 0.114 1.624 0.570 1.609 0.488
 5 Conclusions
@@
 The results revealed that the best model configuration incorporated Frequency Separation, achieving favorable performance in terms of RMSE and Log Spectral Distance (LSD) metrics.
-To operationalize the model, we proposed a downscaling procedure for the entire Canadian domain.
-This procedure generates overlapping downscaled patches of size $128\times 128$ and blends them into the larger Canadian region using a linear interpolation approach.
-As future work, we aim to operationalize the model as a product capable of extending high-resolution wind forecast services from the current 48-hour horizon to the 10-day forecast horizon of the Canadian global forecast model.
+In preparation for an operational environment, we trained the AI model on small $128\times 128$ image patches using a downsampling factor of 8 and made inferences over the entire Canadian domain $1288\times 2536$.
+As future work, we aim to develop the model as a product capable of extending high-resolution wind forecast services from the current 48-hour horizon to the 10-day forecast horizon of the Canadian global forecast model.
 Additionally, we plan to investigate the application of recent state-of-the-art weather foundation models [43] for wind downscaling.